# Post-detection inference for sequential changepoint localization

### v3

@@ G.2.5 Comparison of experimental results with ARL and PFA controlling detection algorithms
G.3 Experimental results with Poisson distribution with detection algorithms that control ARL
G.3.1 Known pre- and post-change (Setting I)
[…]
Figure G.6: Setting I (known pre- and post-change Poisson distributions): The first $T-1$ observations are drawn from Pois(1)(1)( 1 ) and the rest from Pois(2)(2)( 2 ), with $T=100,500$. The point estimate is shown in a vertical red dashed line, and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of 555 independent simulations are shown.
We report the average size and the conditional and unconditional coverage rates across 500500500 independent runs in Table G.7.
Table G.7: Setting I: Average (of 500500500 independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) of the detection algorithm used (LABEL:eq:cusum) with threshold $A$, while varying $T$ and $A$.
$T$ $A$
@@ G.3.1 Known pre- and post-change (Setting I)
G.3.2 Known pre-change and composite post-change (Setting II)
[…]
$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (G.6)
[…]
We report the average size and the conditional and unconditional coverage across 500500500 independent runs in (G.8). The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7). We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
[…]
@@ G.3.2 Known pre-change and composite post-change (Setting II)
G.3.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively. For change detection, we employ a weighted CUSUM-type detector, where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, which is defined as
$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (G.7)
where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section G.3.2 is considered.
After detection, we construct confidence sets (as defined in Section C.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$. We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across 500500500 independent runs in Table G.9.
[…]
@@ G.3.3 Composite pre- and post-change (Setting III)
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table G.7). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table G.7), exhibiting both higher coverage and larger size.
Appendix H A heuristic method for composite pre- or post-change
@@
H.1 Known pre-change and unknown post-change distributions
@@
12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
@@
H.2 Composite pre- and post-change distributions
@@
$\mathcal{H}_{0,t}:X_{1},\cdots,X_{t-1}\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{0},\text{ for some }F_{0}\in\mathcal{P}_{0}\text{ and }X_{t},X_{t+1},\cdots\stackrel{{\scriptstyle i.i.d}}{{\sim}}F_{1},\text{ for some }F_{1}\in\mathcal{P}_{1}.$
@@
As in Algorithm 4, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions. We first obtain some point estimate $\hat{T}$ of the changepoint $T$. Then we use the observations from 111 to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. For some $N\in\mathbb{N}$, we generate $N$ many i.i.d. sequences under $\hat{F}_{0}$ until a change is detected and denote the detection times as $\tau_{1},\cdots,\tau_{N}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with the estimated pre- and post-change distributions $\hat{F}_{0}$ and $\hat{F}_{1}$ respectively, having changepoint at $t$. For each sampled stream $j$, we calculate the test statistic $M_{t}^{j}$ to obtain $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if we have
@@
13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\leavevmode\nobreak\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
@@
H.3 Experimental Results
H.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (B.11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 4 to construct confidence sets. Table H.1 summarizes the results, including the average size and the conditional and unconditional coverage across 500500500 independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table H.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.
@@ H.3.1 Known pre-change and composite post-change (Setting II)
[…]
@@
[…]
H.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and 500500500. We use a weighted CUSUM detector with RIPr defined in (C.12) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. We report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table H.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
Table H.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
@@ H.3.2 Composite pre- and post-change (Setting III)
[…]
@@
[…]
Appendix I Another heuristic method for composite pre- or post-change
@@
For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section H.3.1 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1. Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section H.3.2 and we report the average size and the conditional and unconditional coverage across 500500500 independent runs in Table I.1. From Table I.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
Table I.1: Setting II: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (B.11) with threshold $A=1000$.
@@
[…]
@@
[…]
Table I.2: Setting III: Average (of 500500500 independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (C.12) with threshold $A=1000$.
@@
[…]
@@
[…]

### v4

@@ H.3.4 Comparison of experimental results with ARL and PFA controlling detection algorithms
H.4 Experimental results with Poisson distribution with detection algorithms that control ARL
H.4.1 Known pre- and post-change (Setting I)
We consider Poisson the mean-change problem: $\text{Pois}(1)$ (pre-change) vs $\text{Pois}(2)$ (post-change) with the true (unknown) changepoint at $T=100,500$. We employ the likelihood-based CUSUM change detection method with thresholds at $A=1000,10000$. After detection, we construct confidence sets (adaptive, Equation 4.1) for $T$. Fig. H.6 provides visualizations of the confidence sets and point estimates of the changepoints across $5$ runs.
Figure H.6: Setting I (known pre- and post-change Poisson distributions): The first T1T-1 observations are drawn from Pois(1)(1) and the rest from Pois(2)(2), with $T=100,500$. The point estimate is shown in a vertical red dashed line, and the confidence set (using Algorithm 2) is shown in red points. $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
We report the average size and the conditional and unconditional coverage rates across $500$ independent runs in Table H.7.
Table H.7: Setting I: Average (of $500$ independent runs) size and coverage of the confidence sets using Algorithm 2 with $F_{0}=\text{Pois}(1),F_{1}=\text{Pois}(2),B=N=100$, $\alpha=0.05$, the average absolute deviation $\hat{T}$ and average detection delay (given a true detection) with threshold AA, while varying $T$ and $A$.
$T$ AA
@@ H.4.1 Known pre- and post-change (Setting I)
H.4.2 Known pre-change and composite post-change (Setting II)
We perform experiments for Poisson rate-change scenarios having pre-change distribution as $\text{Pois}(1)$, and the true (unknown) post-change data distribution as $\text{Pois}(2)$, with the changepoint at $T=100,200,$ assuming the composite post-change model to be $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$. For detection, we employ a weighted CUSUM detector defined as
$\tau_{\text{wcusum}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (S5)
for some weight $W(\theta)$ such that $\int_{\theta\in\Theta_{1}}dW(\theta)=1$. We use thresholds $A=1000,10000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, and $\theta=1.9$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we construct confidence sets (as defined in Section C.2), with $\alpha=0.05,\beta=0.025,B=N=100,L=\infty$.
We report the average size and the conditional and unconditional coverage across $500$ independent runs in (H.8). The empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta$ (which is 0.925 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7). We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection) in the table. We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Setting 1), exhibiting both higher coverage and larger size.
Table H.8: Setting II: Average (of $500$ independent runs) size and (conditional and unconditional) coverage rates of confidence sets Section 3.2. with $B=N=100$ and $\alpha=0.05,\beta=0.025,L=\infty$, the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method used (S11) with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
@@ H.4.2 Known pre-change and composite post-change (Setting II)
H.4.3 Composite pre- and post-change (Setting III)
Now we conduct experiments for Gaussian mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$, the true (unknown) pre- and post-change data distributions as $N(0,1)$ and $N(1,1)$ respectively, assuming the composite pre and post-change models to be $\mathcal{P}_{0}=\{\text{Pois}(\theta):\theta<0.9\}$ and $\mathcal{P}_{1}=\{\text{Pois}(\theta):\theta>1.9\}$ respectively. For change detection, we employ a weighted CUSUM-type detector, where the “closest” element from the pre-change (the Reverse Information Projection (RIPr) of the post-change element to the pre-change class) is used as a representative of the pre-change model, which is defined as
$\tau_{\text{wcs-ripr}}=\inf\left\{\tau\in\mathbb{N}:\max_{1\leq j\leq\tau}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{\tau}\frac{f_{\theta}(X_{i})}{f_{\theta}^{*}(X_{i})}dW(\theta)\geq A\right\},$ (S6)
where $f_{\theta}^{*}$ is the RIPr of $f_{\theta}$ on $\mathcal{P}_{0}$. For the post-change parameter, the same discrete weight distribution as in Section H.4.2 is considered.
After detection, we construct confidence sets (as defined in Section D.2, assuming $T\neq 1$), with $\alpha=0.05,\beta=\gamma=0.025,B=N=100,L=\infty$. We report the average size, conditional and unconditional coverage rates, average absolute deviation of $\hat{T}$, and delay (given no false detection) across $500$ independent runs in Table H.9.
Table H.9: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets with $N=B=100$ and $\alpha=0.05,\beta=\gamma=0.025$,the average absolute deviation $\hat{T}$ and average detection delay (given no false detection) of the change detection method with threshold AA, having pre- and post-change data-generating distributions as $\text{Pois}(1)$ and $\text{Pois}(2)$ respectively, while we vary the changepoint $T$ and $A$.
@@ H.4.3 Composite pre- and post-change (Setting III)
However, the empirical conditional coverage rates are always a bit higher than the theoretical bound $1-\alpha-\beta-\gamma$ (which is 0.9 in our experiments). It is worth noting that both the lengths and the conditional coverage rates are higher compared to the known pre- and post-change settings (Table H.7). We observe that the confidence sets in this setting tend to be a bit more conservative compared to those with known pre- and post-change distributions (Table H.7), exhibiting both higher coverage and larger size.
Appendix I A heuristic method for composite pre- or post-change
@@
I.1 Known pre-change and unknown post-change distributions
@@
12:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}F_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}^{\prime}$ using $\mathcal{A}$
@@
I.2 Composite pre- and post-change distributions
@@
[…]
@@
As in Algorithm 4, we employ a simple bootstrap-based approach, but here we need to estimate both pre- and post-change distributions. We first obtain some point estimate $\hat{T}$ of the changepoint $T$. Then we use the observations from $1$ to $\hat{T}-1$ to estimate the pre-change distribution $F_{0}$, denoted $\hat{F}_{0}$ and from $\hat{T}$ to $\tau$ to estimate the post-change distribution $F_{1}$, denoted $\hat{F}_{1}$. For some $N\in\mathbb{N}$, we generate $N$ many i.i.d. sequences under $\hat{F}_{0}$ until a change is detected and denote the detection times as $\tau_{1},\cdots,\tau_{N}$. Then, for some $B\in\mathbb{N}$, we generate $B$ independent bootstrap data streams with the estimated pre- and post-change distributions $\hat{F}_{0}$ and $\hat{F}_{1}$ respectively, having changepoint at $t$. For each sampled stream $j$, we calculate the test statistic $M_{t}^{j}$ to obtain $M_{t}^{1},M_{t}^{2},\dots,M_{t}^{B}$. We reject $\mathcal{H}_{0,t}$ if we have
@@
13:   Simulate $X_{1}^{j},\cdots,X_{t-1}^{j}\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{0};\penalty 10000\ X_{t}^{j},X_{t+1}^{j},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}\hat{F}_{1}$, until a change is detected at $\tau_{j}$ using $\mathcal{A}$
@@
I.3 Experimental Results
I.3.1 Known pre-change and composite post-change (Setting II)
We conduct experiments for both Normal and Poisson mean-change scenarios with a composite post-change model having the changepoint at $T=100$. For detection, we employ a weighted CUSUM detector (S11) with a threshold at $A=1000$ and a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=d=0.2$ and for Poisson distribution $\theta=1.2,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 4 to construct confidence sets. Table I.1 summarizes the results, including the average size and the conditional and unconditional coverage across $500$ independent runs. For $T=100$, conditional and unconditional coverages are equal, as the probability of false detection before the true change is negligible. However, for $T=500$ with $A=1000$, the presence of a significant number of false detections before the true change results in lower unconditional coverage. Nevertheless, the conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this case.
Table I.1: Setting II: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 4 with $B=100,N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (S11) with threshold $A=1000$.
@@ I.3.1 Known pre-change and composite post-change (Setting II)
100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.95 20.95
@@
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1\}$ $\text{Pois}(2)$ 0.95 0.89 20.83
I.3.2 Composite pre- and post-change (Setting III)
Now we conduct experiments for both Normal and Poisson mean-change scenarios with composite pre- and post-change models having the changepoint at $T=100$ and $500$. We use a weighted CUSUM detector with RIPr defined in (S12) for change detection. The thresholds are set at $A=1000$ for both $T=100$ and $T=500$. We considered the same discrete weight distribution as above, taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$ (for normal distribution, we use $\theta=0.8,d=0.2$ and for Poisson distribution $\theta=1.8,d=0.2$) with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized exponential mixture over the alternative class. After detection, we employ Algorithm 5 to construct confidence sets. We report the average size and the conditional and unconditional coverage across $500$ independent runs in Table I.2. The conditional coverage is maintained at the desired level, even though we do not have a theoretical guarantee in this setting.
Table I.2: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 5 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (S12) with threshold $A=1000$.
@@ I.3.2 Composite pre- and post-change (Setting III)
100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.94 0.94 23.32
@@
500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 23.04
Appendix J Another heuristic method for composite pre- or post-change
@@
For the known pre-change and composite post-change setting, we use the same detection algorithm as described in Section I.3.1 and we report the average size and the conditional and unconditional coverage across $500$ independent runs in Table J.1. Similarly, for the composite pre- and post-change setting, we use the same detection algorithm as described in Section I.3.2 and we report the average size and the conditional and unconditional coverage across $500$ independent runs in Table J.1. From Table J.1, we observe that the confidence sets are more conservative when the data-generating distributions are away from those boundaries of the respective composite models.
Table J.1: Setting II: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $B=N=100$ and $\alpha=0.05$, using weighted CUSUM change detection method (S11) with threshold $A=1000$.
@@
100 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.94 0.94 25.83
@@
500 $\text{Pois}(1)$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(2)$ 0.95 0.90 27.17
Table J.2: Setting III: Average (of $500$ independent runs) conditional and unconditional coverage and size of confidence sets using Algorithm 2 with $N=B=100$ and $\alpha=0.05$, using weighted CUSUM type change detection method (S12) with threshold $A=1000$.
@@
100 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.96 0.96 29.64
@@
500 $\{\text{Pois}(\lambda):\lambda\leq 1.25\}$ $\{\text{Pois}(\lambda):\lambda>1.75\}$ $\text{Pois}(1)$ $\text{Pois}(2)$ 0.95 0.95 30.49